ABC is a triangle. If D is the midpoint of AC, show that BA + BC = 2BD.



1
Expert's answer
2021-02-02T03:55:03-0500

Let AB=-x and BC=y , Then AC=AB+BC which is equal to -x+y Then BD=BA+AD =x+1/2(-x+y)=1/2(x+y) which implies that BA+BC=2BD= x+y= 2((1/2(x+y)) hence proved